OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 0 to x = 2.

OpenStudy (anonymous):

x = −4 x = −2 x = 2 x = 4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@thomaster

ganeshie8 (ganeshie8):

average rate of change for f(x) from x = 0 to x = 2 is : \[\large \dfrac{f(2) - f(0)}{2-0}\]

ganeshie8 (ganeshie8):

look at the graph, whats the value of graph at x = 2 ?

OpenStudy (anonymous):

i believe 6 am i right?

OpenStudy (anonymous):

oh ok i see where you are going thanks

ganeshie8 (ganeshie8):

Yep ! \[\large \dfrac{f(2) - f(0)}{2-0}\] \[\large \dfrac{6 - (-2)}{2-0}\] simplify

OpenStudy (anonymous):

you were great help